OpenStudy (anonymous):

what are the coordinates of the terminal point corresponding to θ=15pi/6

OpenStudy (unklerhaukus):

coordinates in x-y or in r-θ ?

OpenStudy (anonymous):

its based on the unit circle

OpenStudy (unklerhaukus):

15/6=2.5 so 15/6 π = 2.5 π = 0.5 π =π/2 the point can be represented as (r, θ) = (1, π/2)

OpenStudy (unklerhaukus):

if you with to represent in x-y use r =√(x^2+y^2) θ = atan2 (y, x)

OpenStudy (anonymous):

thank you so much, i really appreciate it